OpenStudy (anonymous):

I'm new to trigonometry. Can someone help me. Convert each angle in degrees to radians. Express your answer as a multiple of π. 240°.

OpenStudy (lgbasallote):

to change degree to radian just multiply by \(\frac{\pi}{180}\)

OpenStudy (cwrw238):

hint: 180 degrees = pi radians

hero (hero):

Whatever the angle is, multiply it by \(\frac{\pi}{180}\)

hero (hero):

then simplify it

OpenStudy (cwrw238):